OpenStudy (haaileyyy):

Is this Correct? (-5ab+a+5)-(-4ab+5) = -1ab+a

OpenStudy (mariamikayla):

(-5ab+a+5)-(-4ab+5)= = -5ab+a+5 +4ab -5= = -ab +a -the final answer -5 and 5 get reduced

OpenStudy (haaileyyy):

Thank You!

OpenStudy (mariamikayla):

you're welcome :D